how would i solve this quadratic equation? { { { 4x^2=48 } } }
Answer by user_dude2008(1862)

4x^2=48

x^2=48/4

x^2=12

x=+-sqrt(12)

x=2*sqrt(3) or x=-2*sqrt(3) <--------- Answers